OpenStudy (anonymous):

(4-2^2+5)w=25 please show work.

OpenStudy (anonymous):

[4-(2^2)+5)]w = 25 [4-4+5]w=25 (5)w=25 devide both sides by 5 w = 5 :-)

OpenStudy (anonymous):

thank-you so much.

OpenStudy (anonymous):

your welcome! :-)

OpenStudy (anonymous):

also , do you know 7+x-(3+32/8)=3